According to a dietary study, a high sodium intake may be related to ulcers, stomach cancer, and migraine



headaches. The human requirement for salt is only 220 milligrams per day, which is surpassed in most single



servings of ready-to-eat cereals. If a random sample of 15 similar servings of certain cereal has a mean sodium



content of 244 milligrams and a standard deviation of 24.5 milligrams, does this suggest at the 0.05 level of



significance that the average sodium content for a single serving of such cereal is greater than 220 milligrams?

1
Expert's answer
2022-01-17T16:18:58-0500

"H_0: \mu = 220 \\

H_1: \mu > 220 \\

\bar{x} = 244 \\

s = 24.5 \\

n=20 \\

α=0.05 \\

df = n-1 = 20 -1=19"

Test-statistic

"t = \frac{\bar{x} - \mu}{s / \sqrt{n}} \\

t = \frac{244-220}{24.5 / \sqrt{20}} = 4.381 \\

t_{crit} = t_{0.05,19} = 1.729"

Reject H0 if test statistic t>1.729.

Since "t>t_{crit}" , we reject the null hypothesis. We have sufficient evidence to conclude, that "\mu" is greater than 220 at 5% significance level.